Circular No. 7987

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

SUPERNOVA 2002ge IN NGC 7400 R. Chassagne, Ste. Clotilde, Île de Réunion, reports the discovery of

an apparent supernova (mag 14.1) on an unfiltered CCD frame taken with a 0.30-m automated telescope on Oct. 7.90 UT. The new object is located at α = 22^{h}54^{m}21^{s}.49, δ = −45^{o}20′26′′.9 (equinox 2000.0), which is 4′′.0 east and 22′′.6 north of the center of NGC 7400. SN 2002ge was also present on unfiltered CCD frames taken on Oct. 8.63, but it does not appear on unfiltered images taken by Chassagne on 2000 Oct. 1.82 (limiting mag ∼ 18) or on a U.K. Schmidt plate taken on 1976 Sept. 27 (limiting mag 21.0).

SUPERNOVA 2002gd IN NGC 7537 M. Hamuy, Carnegie Observatories, reports that a spectrum (range

380–920 nm) of SN 2002gd (cf. IAUC 7986), obtained on Oct. 8.26 UT with the Las Campanas Dupont 2.5-m telescope (+ WFCCD), shows it to be a type-II supernova with weak Hα, Hβ, and He I 587.6-nm lines exhibiting P-Cyg profiles. The weakness of these lines and the blue continuum suggest a young supernova. The absorption components yield a low expansion velocity of 5300 km/s, based on a redshift of 2560 km/s derived from the narrow Balmer emission lines of the host galaxy.

S. Benetti, G. Altavilla, A. Pastorello, M. Riello, M. Turatto, and L. Zampieri, Osservatorio Astronomico di Padova; E. Cappellaro, Osservatorio Astronomico di Capodimonte; and L. Germany, European Southern Observatory (ESO), on behalf of the European Research and Training Network on the Physics of Type-Ia Supernovae, report that a fully reduced CCD spectrum (range 333–1000 nm, resolution 1.2 nm) of SN 2002gd, obtained on Oct. 8.02 UT with the ESO 3.6-m telescope (+ EFOSC2), shows it to be a type-II supernova caught soon after explosion. The spectrum is dominated by a very blue continuum (T_{bb} ∼ 13000 K) with overimposed hydrogen Balmer and He I 587.6-nm lines showing P-Cyg profiles. From the minima of the lines, an ejecta expansion velocity of ∼ 5000 km/s is derived. The low velocity and the low intrinsic luminosity at such early phases are reminiscent of the peculiar, low-luminosity type-II supernovae (Zampieri et al. 2002, MNRAS, in press).

A. V. Filippenko and R. Chornock, University of California, Berkeley, also write that CCD spectra (range 310–930 nm) obtained on Oct. 8 UT with the Keck-I 10-m telescope (+ Low Resolution Imaging Spectrometer) reveal that SN 2002gd is a young type-II supernova; low-contrast hydrogen Balmer lines with P-Cyg profiles are present on a very blue continuum.

2002 October 8  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2002 CBAT  Daniel W. E. Green